OpenStudy (anonymous):

If f(x) = x + 3 and g(x) = −4x − 1, find g[f(4)]

OpenStudy (anonymous):

for these problems, if you haven't caught on yet you have to plug the 4 into the question of f then into the equation of g \[f(4)=4+3=7, g(f(4))=-7(4)-1=-29\]

OpenStudy (anonymous):

-29 est le anser

hero (hero):

f(x) = x + 3 g(x) = −4x − 1 g(f(x)) = -4(f(x)) - 1 = -4(x + 3) - 1 = -4x - 12 - 1 = -4x - 13 g(f(4)) = -4(4) - 13 = -16 - 13 = ?

OpenStudy (anonymous):

marvin gimme my medal back!!

OpenStudy (anonymous):

you have to plug four into f and u get seven and then plug 7 into g

hero (hero):

For these problems, find g(f(x)) first. Then plug in x = 4 to get g(f(4))

OpenStudy (anonymous):

ur not helping

hero (hero):

@zachy1234, You giving answers is what's not helping at all.

OpenStudy (anonymous):

thank you everyone since i can't give medals to all I'll just fan every1

OpenStudy (anonymous):

@Hero can you explain more